You are performing a right-tailed t - test with test statistic t=0.26t=0.26 and a sample of size 27, find the p-value to 4 decimal places
Answer #1

Solution :

This is the right tailed test .

df = 27 - 1 = 26

P-value = P(t 26 > 0.26) = 0.3985